OpenStudy (anonymous):

Solve the system. y=x+1 y=-3x+13 x= y=

OpenStudy (anonymous):

x = 3 y = 4

OpenStudy (anonymous):

As both equations are equal to y. Equate them both. x + 1 = 13 - 3x 4x = 12 x = 3 Put x into any equation for y. e.g. y = x + 1 -> (3) + 1 = 4 y=4